Show that Sigma infinity n=1 ( -1)^n+1 5÷7n+2 is conditionally convergent

1
Expert's answer
2022-02-22T16:18:37-0500

Solution:

"\sum _{n=1}^{\infty \:}\left(-1\right)^{n+1}\frac{5}{7n+2}
\\=5\cdot \sum _{n=1}^{\infty \:}\left(-1\right)^{n+1}\frac{1}{7n+2}
\\=5\cdot \sum _{n=1}^{\infty \:}\left(-1\right)^n\left(-1\right)\frac{1}{7n+2}
\\=5\left(-\sum _{n=1}^{\infty \:}\left(-1\right)^n\frac{1}{7n+2}\right)
\\=5\left(-\mathrm{converges}\right)\ \ [\text{using Alternating series test}]
\\=\mathrm{converges}"

Here, "a_n=\dfrac1{7n+1}" which is positive and decreasing sequence.

Also, "\lim_{n\rightarrow \infty}a_n=\lim_{n\rightarrow \infty}\dfrac1{7n+1}=0"

Thus, by alternating series test "\sum _{n=1}^{\infty \:}\left(-1\right)^{n}a_n\ or\ \sum _{n=1}^{\infty \:}\left(-1\right)^{n+1}a_n" converge.